Im thinking of a whole number. If I divide it by 12, then multiply the answer by7, then subtract 5, and then add 13, I end up with 15. What was my original number?
Found 2 solutions by ewatrrr, ptaylor:
Answer by ewatrrr(24785)


Hi
Let x represent the number
Write as we Read***
(x/12)*7 - 5 + 13 = 15
Solving for x
7x - 60 + 13*12 = 15*12 Multiplying @ term by 12 so as all denominaors = 1
7x - 60 + 156 = 180
7x = 84
x = 12
CHECKING our Answer***
7 - 5 + 13 = 15

Answer by ptaylor(2198)

Let x=the whole #
Divide it by 12: x/12
Multiply ans by 7: 7(x/12)=7x/12
Then subtract 5: (7x/12)-5
And then add 13: ((7x/12)-5)+13
And the above equals 15: Sooooo
((7x/12)-5)+13=15 get rid of parens
7x/12-5+13=15 simplify
7x/12 +8=15 subtract 8 from each side
7x/12=7 multiply each side by 12
7x=84
x=12----the whole #
CK
divide by 12=12/12=1
Multiply by 7=7*1=7
Subtract 5=7-5=2
Add 13=13+2=15
15=15
Hope this helps-ptaylor